# DynamicPAE: Generating Scene-Aware Physical Adversarial Examples in Real-Time

### v1

@@ Abstract
Physical adversarial examples (PAEs) are regarded as “whistle-blowers” of real-world risks in deep-learning applications, thus worth further investigation. However, current PAE generation studies show limited adaptive attacking ability to diverse and varying scenes, revealing the emergency requirement of dynamic PAEs that are generated in real-time and conditioned on the observation from the attacker. The key challenges in generating dynamic PAEs are exploring their patterns under noisy gradient feedback and adapting the attack to agnostic scenario natures. To address the problems, we present DynamicPAE, the first generative framework that enables scene-aware real-time physical attacks beyond static attacks. Specifically, to train the dynamic PAE generator under noisy gradient feedback, we introduce the residual-driven sample trajectory guidance technique, which is mainly achieved by redefining the training task to break the limited feedback information restriction that leads to the degeneracy problem. Intuitively, it allows the gradient feedback to be passed to the generator through a low-noise auxiliary task, thereby guiding the optimization away from degenerate solutions and facilitating a more comprehensive and stable exploration of feasible PAEs. To adapt the generator to agnostic scenario natures, we introduce the context-aligned scene expectation simulation process, consisting of the conditional-uncertainty-aligned data module and the skewness-aligned objective re-weighting module. The former enhances robustness in the context of incomplete observation by employing a conditional probabilistic model for domain randomization, while the latter facilitates consistent stealth control across different attack targets by automatically reweighting losses based on the skewness indicator. Extensive digital and physical evaluations demonstrate the superior attack performance of DynamicPAE, attaining a 1.95× boost (65.55% average AP drop under attack) on representative object detectors (e.g., Yolo-v8) over state-of-the-art static PAE generating methods. Overall, our work opens the door to end-to-end modeling of dynamic PAEs by overcoming the uncertainty inherent in their generation.
@@ 1 Introduction
Numerous intelligent applications in real-world scenarios have been landed in recent years, such as autonomous driving [1], healthcare [2] and intelligent assistant [3]. However, adversarial examples (AEs), which are specially designed for misleading machine learning models [4], have long been a challenge for deep learning applications [5]. Among them, the physical-world adversarial examples (PAEs) are attracting broader attention due to their feasibility in the real world and de facto threats to business AI systems [6, 7, 8, 9]. Besides revealing the risks, the research on AEs (PAEs) also deepens the understanding of deep neural networks and reveals their defects in industrial and science applications [10, 11]. Therefore, modeling AEs, especially PAEs, is worth further investigating to alleviate the trustworthy issues (e.g., interpretability, security, and robustness) of deep learning models and related applications.
@@
Figure 1: DynamicPAE framework trains real-time scene-ware physical adversarial example (PAE) generators and establishes the novel form of PAE, i.e., dynamic PAE.
@@
The existing adversarial attacks can be categorized as digital attacks and physical attacks, carried out in digital and real-world scenarios, respectively. On this basis, a line of studies aims to further propose adversarial attacks and promote the explainability of deep learning models [12, 13, 14, 15]. Another line of research focuses on realizing PAEs in real-world settings [16, 17], as digital AEs may not be robust when fabricated in the real world [18]. In terms of threat, PAEs pose greater social risks, therefore highly attracting broader focuses. However, mainstream PAE methods address the real-world PAE generation problem by treating it as a static issue, either attempting to generalize the PAE across all simulated physical scenes [19] or requiring retraining each time to adapt to new settings [20]. This approach results in inadequate adaptability and generation efficiency. On the contrary, we define scene-aware and real-time generated PAEs as dynamic PAEs, which is a problem that has yet to be addressed in the field. Related attempts are limited to determining patch locations only [21], simulated control of a few states [22], optimization for each clustered scene in the lab [23], and resisting dynamic fluctuations [24] rather than enabling dynamic responses.
This work focuses on overcoming the core challenge, i.e., the key inherent uncertainty of PAEs, of dynamic PAE generation. It is a new generative task that takes the observation from the attacker as the input and the corresponding PAE as the output. However, compared to conventional generative tasks that learn from existing data, this task requires creating unknown adversarial data that is effective in the uncertain world, and the following challenges stem from modeling the inherent uncertainty of the PAE generation task need to be tackled: ❶ Training the dynamic PAE generator is difficult, presented as optimization degeneracy and infinite gradients. The difficulty arises from the noisy gradient feedback [25] in attack training that hinders the exploration of the potentially sparse distributions [26] of PAEs, and the noise originates from the necessary randomness injected into PAE training. ❷ The real-world attack scenario contains agnostic properties, including incomplete observations that limit the perception of the generator and different attack targets that influence the balance between objectives. Therefore, further modeling of scene-aware attacks is required to eliminate the physical-digital gap and balance the objectives for consistent generative behavior.
@@
To address the dynamic PAE generation problem, we present the DynamicPAE framework that effectively addresses the challenges. As shown in Figure 1, the framework trains the scene-aware generator in the end-to-end paradigm and significantly improves the attack performance under varying conditions. Specifically, to address the training problem that arises from noisy gradients, residual-driven sample trajectory guidance is proposed by breaking the limited feedback information restriction that leads to the degeneracy problem. Inspired by deep residual learning [27], we present the construction of the residual task that redefines the optimization problem, enriches the guidance information by relaxing the training with a low-noise auxiliary task, and leads the optimization toward exploring more aggressive results in the sparse space of PAEs. A regularized latent encoding is further proposed to stabilize the exploration. To adapt the agnostic attack scenarios, context-aligned scene expectation simulation is proposed. Specifically, the conditional-uncertainty-aligned data module is proposed to inject the randomness that is aligned with the real-world uncertainty, including the uncertainty of incomplete observation, and the skewness-based weight controller is proposed to automatically align the objectives when directly adapting the framework to agnostic target models. By solving the key training and task modeling issues, the framework bridges the gap between generative neural networks and dynamic PAE generation.
@@
  • We propose the novel dynamic physical adversarial example (PAE), that is generated conditioned on the observation from the attacker in real-time. We propose the DynamicPAE framework that opens the end-to-end modeling of this widely impactful task.
  • We propose the residual-driven sample trajectory guidance technique, which addresses the degeneracy and instability problem in the generative training of dynamic PAEs. We recognize that the problem arises from the noisy gradient feedback of attack training, and our technique breaks the limited feedback information restriction, which leads to the degeneracy problem, by redefining the training task and allowing the gradient feedback to be passed to the generator through a low-noise auxiliary task. It guides the optimization away from degenerate solutions, leading to a more comprehensive and stable exploration of feasible PAEs.
  • We propose the context-aligned scene expectation simulation process for addressing the agnostic nature of attack scenarios. It consists of the conditional-uncertainty-aligned data module that improves the robustness under incomplete observations by utilizing domain randomization through the conditional probabilistic model, and the skewness-aligned objective re-weighting module that enables consistent stealthiness control across different attack targets by automatically re-weighting losses using the skewness indicator.
  • Extensive experiments across diverse settings demonstrate the superiority of DynamicPAE, i.e., 1.95$\times$ of average AP drop after attack compared to existing PAEs and 12ms inference latency. Evaluations on the simulation platform (i.e., CARLA) and physical-world-deployed devices further validate the effectiveness of our proposed method.
The rest of this paper is organized as follows: section 2 presents the background and the definition of the dynamic PAE problem. Section 3 introduces the key components of the DynamicPAE framework. Section 4 details the experimental evaluation, including the settings, the results, and the analysis. Section 5 discusses the related works of the dynamic attack problem and optimization techniques, and section 6 concludes the paper with possible future works.
@@ 2.1 Backgrounds
In the past decade [33], algorithms supporting PAE generation have mainly focused on applying the iterative optimization with EoT paradigm to find a universal PAE that is robust under all simulated scenes  [34, 35, 20, 36, 24]. On the contrary, the solutions based on the inference of non-iterative models, e.g., conditional generative networks [37], have been validated only in the digital or randomness-free scenarios. However, the iterative paradigm does not support scene-adaptive PAEs because it requires inefficient repetitive optimization [38, 9]. It is difficult to fully achieve the capability to present the corresponding PAE by current observation, i.e., scene-aware, for the attackers.
@@ 2.2 Definitions
[…]
@@
$\mathcal{L}_{\mathrm{GCAtk}};$ Guided and Controlled Attack Loss
@@ 3 Methodology
The overview of the DynamicPAE framework is shown in Figure 2. We aim to establish an end-to-end framework for dynamic PAE generation based on the unified representation learning paradigm, which formulates the architecture of the neural PAE generator as follows.
@@
To enable effective training of the PAE generator $\mathcal{G}$, the framework synergizes two key components: residual-driven sample trajectory guidance to tackle the training difficulties arise from noisy gradient feedback; and context-aligned scene expectation simulation to adapt to agnostic attack scenarios, including incomplete observations and agnostic attack targets, thereby achieving the dynamic adversarial attack in the real world.
Figure 2: The overview of the DynamicPAE  framework proposed for the dynamic physical adversarial attack problem. The framework establishes the training paradigm of Residual-Driven Sample Trajectory Guidance that solves the training difficulties arise from noisy gradient feedback, and the Context-Aligned Scene Expectation Simulation that adapts the model to agnostic scenarios.
3.1 Residual-Driven Sample Trajectory Guidance
Training the dynamic PAE generator faces the challenge of noisy gradient feedback, which leads to the degeneracy and instability problem of the generative training (shown in the experiments in section VII). To analyze the training problem, we bridge the noisy gradient model and the sample trajectory analysis with the limited feedback information restriction and recognize that the necessary randomness injected in the PAE optimization task leads to the training degeneracy of the dynamic PAE generator. Inspired by deep residual learning, we tackle the problem by constructing a residual task that relaxes the original attack training and breaks the limited feedback information restriction. We further propose the additional regularization of the latent encoding to tackle the latent evasion problem related to the unconventional gradient property and stabilize the training. Afterward, the training escapes from degenerate solutions and allows a more stable and comprehensive exploration of feasible PAEs.
$\bullet$ Analysis of the Degeneracy: Previous works have demonstrated that obfuscated gradient [25] is a mechanism of adversarial defense techniques and adaptive attacks shall overcome it [44]. Section 4.3.1 provides compelling evidence that the simulated transformation has a similar impact on PGD optimization, confirming the existence of the noisy gradient problem. Intuitively, a small distortion in the environment $(\mathbf{X},\boldsymbol{\theta})$ may cause the gradient feedback of the attack task $\nabla_{\boldsymbol{\delta}}\mathcal{L}_{\mathrm{Atk}}$ changes significantly. We formulate the limited feedback information restriction to model the problem based on the mutual information and entropy metrics on $\nabla_{\boldsymbol{\delta}}\mathcal{L}_{\mathrm{Atk}}$ and the (bottleneck) latent variable $\mathbf{Z}$:
@@ 3.1 Residual-Driven Sample Trajectory Guidance
where $\eta$ is the learning rate, $g^{(t)}$ is the noise introduced by random simulated transformation, and the term $f^{(t)}$ is the components correlated with $\mathbf{Z}$, representing the learnable information. By rate-distortion theory, the signal power $\mathbb{E}[|f^{(t)}|^{2}]$ is significantly lower than the noise power $\mathbb{E}[|g^{(t)}|^{2}]$ if $I(\nabla_{\boldsymbol{\delta}}\mathcal{L}_{\mathrm{Atk}}^{(t)};\mathbf{Z})$ is relatively small according to Eq. 4. Based on the empirical fact of the existence of universal AE [45], the AE optimization contains isotropic components, and $\mathbb{E}[g^{(t)}]$ is large. Thus the optimization of PAE may converge to the local minimum determined by $g^{(t)}$ and $\mathbb{E}[g^{(t)}]$ even with sufficient large steps $t$, which is coherent with the experiment results in section 4.3.1.
This result indicates the hardness of SGD optimization in learning PAEs and is independent of whether the optimizer includes techniques like momentum to escape from local optima. Generative models, especially single-step models such as GAN, have long faced the similar experimental challenge of mode collapse in the manifold sampling tasks, in particular on disconnected manifolds [46, 47, 48]. Rather than relying on generative model construction techniques, we propose a task-driven approach and formulate solutions that are well-suited to the problem itself.
$\bullet$ Optimization Guidance with Residual Task: Inspired by deep residual learning [27], we bypass this difficulty by redefining the optimization task. Specifically, we simply relax the adversarial attack task with the additional task $\mathcal{R}$ with the conditional parameter $\lambda$, and modify the loss from $\mathcal{L}_{\mathrm{Atk}}(\boldsymbol{\delta}(\cdot))$ to $\mathscr{L}_{\lambda}(\mathcal{L}_{\mathrm{Atk}},\mathcal{L}_{\mathcal{R}})$, defined as:
$\mathscr{L}_{\lambda}(\mathcal{L}_{\mathrm{Atk}},\mathcal{L}_{\mathcal{R}}):=\lambda\mathcal{L}_{\mathrm{Atk}}(\boldsymbol{\delta}(\cdot,\lambda))+(1-\lambda)\mathcal{L}_{\mathcal{R}}(\boldsymbol{\delta}(\cdot,\lambda)),$ (6)
@@
$\exists\lambda\text{ s.t. }I(\nabla_{\boldsymbol{\delta}}\mathscr{L}_{\lambda}(\mathcal{L}_{\mathrm{Atk}},\mathcal{L}_{\mathcal{R}});\mathbf{Z}(\cdot,\lambda))\gg I(\nabla_{\boldsymbol{\delta}}\mathcal{L}_{\mathrm{Atk}};\mathbf{Z}).$ (7)
Afterward, the imbalanced noise issue in SGD optimization is solved, which mitigates the occurrence of degeneration. Further, the optimal PAE solution remains identical when $\lambda=0$. Since the gradient takes a noiseless path, or a residual path, to the parameters of $\mathcal{G}$, we identify task $\mathcal{R}$ as the auxiliary residual task.
More intuitively, the goal of the residual task in generating PAEs is to encourage the exploration of the global space of PAEs. Inspired by the construction of the denoising task in diffusion models to learn the entire gradient field [49], we designate local area reconstruction as the auxiliary residual task $\mathcal{R}$ ($\mathcal{L}_{\mathcal{R}}:=\mathcal{L}_{\mathrm{Inv}}$) since it is directly related to the model input, is well studied by generative models, and lets the model learns to generate PAEs at different magnitudes, possessing the utility of controllable stealthiness. Specifically, we adapt the mean square error (MSE) as the objective quality metric and the LPIPS [50] as the subjective quality metric, and integrate them as:
@@
where $\mathbf{X}$ and $\mathbf{X}^{\prime}=\mathbf{X}\oplus\boldsymbol{\delta}$ is the global image capture before and after applying PAE $\boldsymbol{\delta}$, respectively.
@@
where $\mathrm{\delta}$ denotes the Dirac distribution and $b$ denotes the batch size. The distribution is constructed by sampling minimal, maximal, and random hyper-parameters and is inspired by the sampling techniques in auto-ML training [51]. To generate PAEs conditioned on $\lambda$, $\lambda$ is encoded into the latent representation $\mathbf{Z}$ in $\mathcal{G}$ by $\log$ embeddings and a $\mathrm{MLP}$ module modeling the non-linear transformation:
@@
$\mathbf{Z}^{\prime}=\mathrm{MLP}(-\log(\max\{\boldsymbol{\lambda},\exp(-10)\})+\mathbf{Z}.$ (10)
@@
Afterward, the generator $\mathcal{G}$ generates PAEs with conditioned magnitude $\boldsymbol{\delta}=[\boldsymbol{\delta}_{\lambda_{1}},\boldsymbol{\delta}_{\lambda_{2}},...,\boldsymbol{\delta}_{\lambda_{b}}]$ in each batch, and the same $\boldsymbol{\lambda}$ is applied to Eq. 6 as training objectives.
@@
$\bullet$ Stabilizing Sample Exploration: The particularity of the dynamic PAE generation also exists in the latent space of generator $\mathcal{G}$. Empirically, we observed a non-trivial gradient explosion behavior in the generator during training. The analysis is shown in section 4.3.2 and the supplementary material, indicating the difference between PAE generation and conventional tasks. We recognize it as the latent evasion problem, which stems from the process of exploring PAEs.
@@
Specifically, due to the analyzed noisy gradient problem, the discovery of new conditional PAEs often requires multiple steps of fitting to emerge. Meanwhile, the disentangled representations and weights within the model continuously deviate towards a certain direction throughout the multi-step fitting process, causing their norms to continually increase. Furthermore, the overall discontinuity of the PAE distribution further intensifies this phenomenon. Therefore, the optimization speed of the model’s internal representations should be slowed down. Thus, to tackle this problem, we incorporate a regularization loss on the latent representation of PAEs $\mathbf{Z}$ to stabilize the exploration, which together regularizes the PAEs with the total variation loss $\mathcal{L}_{\mathrm{TV}}$ [52] on the data space that eliminates the frequency discrepancy between PAEs and physical realizable images:
$\mathcal{L}_{\mathrm{Reg}}:=\gamma\cdot||\mathbf{Z}||_{2}^{2}+\beta\cdot\mathcal{L}_{\mathrm{TV}},$ (11)

### v3

@@ Abstract
Physical adversarial examples (PAEs) are regarded as “whistle-blowers” of real-world risks in deep-learning applications, thus worth further investigation. However, current PAE generation studies show limited adaptive attacking ability to diverse and varying scenes, revealing the urgent requirement of dynamic PAEs that are generated in real time and conditioned on the observation from the attacker. The key challenge in generating dynamic PAEs is learning the sparse relation between PAEs and the observation of attackers under the noisy feedback of attack training. To address the challenge, we present DynamicPAE, the first generative framework that enables scene-aware real-time physical attacks. Specifically, to address the noisy feedback problem that obfuscates the exploration of scene-related PAEs, we introduce the residual-guided adversarial pattern exploration technique. We first introduce the limited feedback information restriction to model the training degeneracy problem under noisy feedback. Then, residual-guided training, which relaxes the attack training with a reconstruction task, is proposed to enrich the feedback information, thereby achieving a more comprehensive exploration of PAEs. To address the alignment problem between the trained generator, which represents the learned relation, and the real-world scenario, we introduce the distribution-matched attack scenario alignment, consisting of the conditional-uncertainty-aligned data module and the skewness-aligned objective re-weighting module. The former aligns the training environment with the incomplete observation of the real-world attacker. The latter facilitates consistent stealth control across different attack targets by balancing the objectives with the skewness indicator. Extensive digital and physical evaluations demonstrate the superior attack performance of DynamicPAE, attaining a 2.07× boost (58.8% average AP drop under attack) on representative object detectors (e.g., DETR) over state-of-the-art static PAE generating methods. Overall, our work opens the door to end-to-end modeling of dynamic PAEs.
@@
©2025 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works. DOI: 10.1109/TPAMI.2025.3626068
@@ 1 Introduction
Numerous intelligent applications in real-world scenarios have been landed in recent years, such as autonomous driving [1], healthcare [2] and intelligent assistant [3]. However, adversarial examples (AEs), which are specially designed for misleading machine learning models [4], have long been a challenge for deep learning applications [5]. Among them, the physical-world adversarial examples (PAEs) are attracting broader attention due to their feasibility in the real world and de facto threats to business AI systems [6, 7, 8, 9]. Besides revealing the risks, the research on AEs (PAEs) also deepens the understanding of deep neural networks and reveals their defects in industrial and scientific applications [10, 11]. Therefore, modeling AEs, especially PAEs, is worth further investigation to alleviate the trustworthy issues (e.g., interpretability, security, and robustness) of deep learning models and related applications.
@@
Figure 1: DynamicPAE framework trains real-time scene-aware physical adversarial example (PAE) generators and establishes a novel form of PAE, i.e., dynamic PAE. After training, the generator generates highly effective PAEs based on the characteristics of the current scene.
@@
Generally, PAEs might target varying real-world conditions in application scenarios, such as moving objects, dynamic lighting, and ever-changing backgrounds. However, mainstream PAE methods address the PAE generation problem by treating it as a static issue, either attempting to generalize the PAE across all simulated physical scenes [12] or requiring retraining each time to adapt to new settings [13]. This approach results in inadequate adaptability or generation efficiency. On the contrary, we define scene-aware and real-time generated PAEs as dynamic PAEs, a problem that has yet to be addressed in the field. Related attempts are limited to determining patch locations only [14], simulated control of a few states [15], optimization for each clustered scene in the lab [16], and resisting dynamic fluctuations [17] rather than enabling dynamic responses.
This work focuses on constructing the dynamic PAE generating framework to establish the novel form of PAE, i.e., dynamic PAE.x As shown in Figure 1, we aim to construct the real-time dynamic PAE generator through an end-to-end training and deployment paradigm, achieving highly effective PAE generation by adapting to the current scene in real time. This can be regarded as a new generative learning problem that takes the observation from the attacker as the input and the corresponding PAE as the output. However, compared to conventional generative learning tasks that extract knowledge from existing data, this task requires creating unknown adversarial data that is effective in an uncertain world, and the generator shall learn the key relations between the observation of the attacker and suitable PAEs. Such relation may be sparse, and the optimization feedback of the attack task may be noisy, which poses the following challenges: ❶ The suitable scene-related PAEs are hard to discover, presented as the optimization degeneracy problem in training dynamic PAE generators. The problem arises from the noisy gradient feedback [18] in attack training that hinders the exploration of the potentially sparse distributions [19] of PAEs, and the noise originates from the necessary randomness injection simulating the real-world uncertainty. ❷ Aligning the trained generator, which represents the learned relations, with the real-world scenario is challenging. The observation of the real-world attacker is incomplete, and the behavior of victim models is inconsistent. These characteristics can cause difficulties in balancing universality, stealth, and attack performance when adapting the generator to real-world scenarios.
@@
To address the dynamic PAE generation problem, we present the DynamicPAE framework that guides the design and the training of the dynamic PAE generator. As shown in Figure 1, the framework trains the scene-aware generator in the end-to-end paradigm and significantly improves the attack performance under varying conditions. Specifically, residual-guided adversarial pattern exploration is proposed to address the training degeneracy problem of the dynamic PAE generator. We first propose the limited feedback information restriction model to analyze the training degeneracy problem and support the construction of training techniques. Inspired by deep residual learning [20], we present the construction of the residual task that redefines the optimization problem, enriches the guidance information by relaxing the training with an auxiliary task, and leads the optimization toward the deep exploration of the relation between PAEs and scenes. Distribution-matched attack scenario alignment is proposed to align the generator with the real-world scenario. Specifically, the conditional-uncertainty-aligned data model is proposed to align the training environment with the incomplete observation of the attacker and therefore balance the trade-off between the universality and attack performance. The skewness-based weight controller is proposed to align the stealthiness of generated PAEs automatically by the novel skewness indicator, achieving the consistent and balanced generating objective control across different attack targets. By solving the key training and task modeling issues, the DynamicPAE framework bridges the gap between generative neural networks and dynamic PAE generation.
@@
  • We propose the novel dynamic physical adversarial example (PAE), which is generated conditioned on the observation from the attacker in real-time. We propose the DynamicPAE framework that opens the end-to-end modeling of this widely impactful task.
  • We propose the residual-guided adversarial pattern exploration to address the training degeneracy problem and explore suitable PAEs under noisy feedback. We model the problem by the limited feedback information restriction. We propose the residual-guided training technique that relax the attack training with a reconstruction task to enrich the feedback information. It guides the optimization away from degenerate solutions, leading to a more comprehensive and stable exploration of feasible PAEs.
  • We propose the distribution-matched attack scenario alignment to address the challenges of aligning the generator with the real-world scenario. It consists of (1) the conditional-uncertainty-aligned data module that models the incomplete observation of the attacker and aligns the training environment with it; (2) the skewness-aligned objective re-weighting module that enables consistent stealthiness control across different attack targets by automatically re-weighting losses using the skewness indicator.
[…]
The rest of this paper is organized as follows: Section 2 presents the background and the definition of the dynamic PAE problem. Section 3 introduces the key components of the DynamicPAE framework. Section 4 details the experimental evaluation, including the settings, the results, and the analysis. Section 5 discusses the related works of the dynamic attack problem and optimization techniques. Section 6 discusses the potential application, limitation and future works, and section 7 concludes the paper.
@@ 2.1 Backgrounds
In the past decade [27], algorithms supporting PAE generation have mainly focused on applying the iterative optimization with EoT paradigm to find a universal PAE that is robust under all simulated scenes [28, 29, 13, 30, 17]. On the contrary, the solutions based on the inference of non-iterative models, e.g., conditional generative networks [31], have been validated only in the digital or randomness-free scenarios. However, the iterative paradigm does not support scene-adaptive PAEs because it requires inefficient repetitive optimization [32, 9]. It is difficult to fully achieve the capability to present the corresponding PAE by current observation, i.e., scene-aware, for the attackers. We overcome the obstacle by advancing the training technology of efficient generative neural networks for this task.
@@ 2.2 Definitions
$\textit{find }\mathcal{G}\text{ s.t. }\mathcal{F}(\mathbf{X}\oplus(\mathcal{G}(\mathbf{P}_{\mathrm{X}}),\boldsymbol{\theta}))\in Y_{adv},$ (2)
@@ 3 Methodology
The overview of the DynamicPAE framework is shown in Figure 2. To solve the dynamic PAE generation problem and achieve real-time performance, modeling the relation within the simplified and unified representation is crucial, which formulates the general neural architecture of the PAE generator as follows.
@@
DynamicPAE framework provides the key guidance for the training and the design of $\mathcal{G}$. The framework synergizes two key components: residual-guided adversarial pattern exploration that tackles the training difficulties of $\mathcal{G}$; and distribution-matched attack scenario alignment that adapts $\mathcal{G}$ to the physical world, thereby achieving the scene-aware dynamic adversarial attack in the real world. Based on these components, DynamicPAE establishes an end-to-end training pipline for $\mathcal{G}$.
Figure 2: The overview of the DynamicPAE framework, consisting of Residual Guided Adversarial Pattern Exploration that solves the training difficulties and the design methodology of Distribution-Matched Attack Scenario Alignment that aligns the generator to real-world constraints. The former enriches the guidance information with a residual-constructed reconstruction task, which decouples and diversifies the exploration of scene-conditional PAEs (shown as green & blue arrows). The latter adapts a probabilistic simulation process to model the real-world uncertainty, and a skewness-aligned objective weight controller for automatically balancing the stealthiness and attack objectives.
3.1 Residual-Guided Adversarial Pattern Exploration
We first focus on the non-trivial training problem of the generator $\mathcal{G}$ under Eq. 2. We found that the training degeneracy problem always occurs in this task (shown in the experiments in section IX). To characterize the key difficulty, we propose the limited feedback information restriction model that accounts for the degeneracy of the generative training. Inspired by deep residual learning, we tackle the problem by constructing a residual task that relaxes the original attack training and breaks the limited feedback information restriction. We further propose the regularized latent encoding to tackle the latent evasion problem and achieve stabilized multi-task training. Afterward, the training escapes from degenerate solutions and allows a more stable and comprehensive exploration of feasible scene-related PAEs.
$\bullet$ Characterizing the Degeneracy: We utilize an information-theoretic modeling approach to construct a concise model for the training problem and then establish a connection between noisy gradients and the trajectories of generated samples. Previous works have demonstrated that obfuscated gradient [18] is a mechanism of adversarial defense techniques, and adaptive attacks shall overcome it [39]. Section 4.3 provides compelling evidence that the simulated transformation has a similar impact on PGD optimization, confirming the existence of the noisy gradient problem. Intuitively, a small distortion in the environment $(\mathbf{X},\boldsymbol{\theta})$ may cause the gradient feedback of the attack task $\nabla_{\boldsymbol{\delta}}\mathcal{L}_{\mathrm{Atk}}$ changes significantly. We formulate the limited feedback information restriction to model the problem based on the mutual information and entropy metrics on $\nabla_{\boldsymbol{\delta}}\mathcal{L}_{\mathrm{Atk}}$ and the (bottleneck) latent variable $\mathbf{Z}$:
@@ 3.1 Residual-Guided Adversarial Pattern Exploration
where $\eta$ is the learning rate, $g^{(t)}$ is the noise introduced by random simulated transformation, and the term $f^{(t)}$ is the components correlated with $\mathbf{Z}$, representing the learnable information. By rate-distortion theory, the signal power $\mathbb{E}[|f^{(t)}|^{2}]$ is significantly lower than the noise power $\mathbb{E}[|g^{(t)}|^{2}]$ if $I(\nabla_{\boldsymbol{\delta}}\mathcal{L}_{\mathrm{Atk}}^{(t)};\mathbf{Z})$ is relatively small according to Eq. 4. Based on the empirical fact of the existence of universal AE [40], the AE optimization contains isotropic components, and $\mathbb{E}[g^{(t)}]$ is large. Thus, the optimization of PAE may converge to the local minimum determined by $g^{(t)}$ and $\mathbb{E}[g^{(t)}]$ even with sufficiently large steps $t$, which is coherent with the experiment results in section 4.3.
Figure 3: Training Problem and Solution Illustration.
A more intuitive illustration is shown in Figure 3. Specifically, the original training method without the residual task is unsatisfactory due to the limited feedback information constraint. Our qualitative analysis shows that it leads to the degeneracy of training, i.e. the generator only generates a single form of PAE, which is independent of whether the optimizer includes techniques like momentum to escape from local optima. The advancements in generative models, such as diffusion models [41] and autoregressive models [42], have significantly improved the quality of complex data generation with better training task construction. However, the degeneracy highlights the unique challenges inherent to the task of dynamic PAE training. Therefore, we argue that addressing it requires a fundamental transformation of the training task itself. As a solution, we propose the residual-guided training that breaks the limited feedback information constraint by redefining the training task, leading to the decoupled and diversified exploration of PAEs.
$\bullet$ Exploration Enhancement with Residual Task: As shown in the right part of Figure 3, we bypass the difficulty by redefining the training task to enrich the feedback information. Specifically, we relax the adversarial attack task with the additional task $\mathcal{R}$ with the conditional parameter $\lambda$, and modify the loss from $\mathcal{L}_{\mathrm{Atk}}(\boldsymbol{\delta},\cdot)$ to $\mathscr{L}_{\lambda}(\mathcal{L}_{\mathrm{Atk}},\mathcal{L}_{\mathcal{R}})$, defined as:
$\mathscr{L}_{\lambda}(\mathcal{L}_{\mathrm{Atk}},\mathcal{L}_{\mathcal{R}}):=\lambda\mathcal{L}_{\mathrm{Atk}}(\boldsymbol{\delta}_{\lambda},\cdot)+(1-\lambda)\mathcal{L}_{\mathcal{R}}(\boldsymbol{\delta}_{\lambda},\cdot).$ (6)
@@
$\exists\lambda\text{ s.t. }I(\nabla_{\boldsymbol{\delta}_{\lambda}}\mathscr{L}_{\lambda}(\mathcal{L}_{\mathrm{Atk}},\mathcal{L}_{\mathcal{R}});\mathbf{Z}_{\lambda})\gg I(\nabla_{\boldsymbol{\delta}}\mathcal{L}_{\mathrm{Atk}};\mathbf{Z}).$ (7)
Afterward, the imbalanced noise issue in SGD optimization could be solved, which mitigates the occurrence of degeneration. Further, the optimal PAE solution remains identical when $\lambda=0$. Since the gradient takes a noiseless path, or a residual path, to the parameters of $\mathcal{G}$, we identify task $\mathcal{R}$ as the auxiliary residual task.
From the perspective of the PAE generating space, the goal of the residual task in generating PAEs is to encourage the exploration of the global space of PAEs. Inspired by the construction of the denoising task in diffusion models to learn the entire gradient field [41], we designate local area reconstruction as the auxiliary residual task $\mathcal{R}$ ($\mathcal{L}_{\mathcal{R}}:=\mathcal{L}_{\mathrm{Inv}}$) since it is directly related to the model input $\mathbf{P}_{\mathrm{X}}$, is well studied by generative models, and lets the model learns to generate PAEs at different magnitudes, possessing the utility of controllable stealthiness. Specifically, we adapt the mean square error (MSE) as the objective quality metric and the LPIPS [43] as the subjective quality metric, and integrate them as:
@@
where $\mathbf{X}$ and $\mathbf{X}^{\prime}=\mathbf{X}\oplus\boldsymbol{\delta}$ are the global image captures before and after applying PAE $\boldsymbol{\delta}$, respectively.
@@
where $\mathrm{\delta}$ denotes the Dirac distribution and $b$ denotes the batch size. The distribution is constructed by sampling minimal, maximal, and random hyperparameters and is inspired by the sampling techniques in AutoML training [44]. To generate PAEs conditioned on $\lambda$, $\lambda$ is encoded into the latent representation $\mathbf{Z}$ in $\mathcal{G}$ by $\log$ embeddings and a $\mathrm{MLP}$ module modeling the non-linear transformation:
@@
$\mathbf{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}Z}_{\boldsymbol{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\lambda}}=\mathrm{MLP}(-\log(\max\{\boldsymbol{\lambda},\exp(-10)\})+\mathbf{Z}.$ (10)
@@
We denote the $\boldsymbol{\lambda}$-conditioned generator as $\mathcal{G}_{\boldsymbol{\lambda}}$, which generates PAEs $\boldsymbol{\delta}_{\boldsymbol{\lambda}}=[\boldsymbol{\delta}_{\lambda_{1}},\boldsymbol{\delta}_{\lambda_{2}},...,\boldsymbol{\delta}_{\lambda_{b}}]$ with conditioned magnitudes in each training batch, and the same $\boldsymbol{\lambda}$ is applied to Eq. 6 as training objectives.
@@
$\bullet$ Stabilizing Sample Exploration: The particularity of the dynamic PAE generation also exists in the latent space of generator $\mathcal{G}$. Empirically, we observed a non-trivial gradient explosion behavior in the generator during training. The analysis is shown in the supplementary material, indicating the difference between PAE generation and conventional tasks. We recognize it as the latent evasion problem, which stems from the process of exploring PAEs.
@@
Specifically, due to the analyzed noisy gradient problem, the discovery of new conditional PAEs often requires multiple steps of fitting to emerge. Meanwhile, the disentangled representations and weights within the model continuously deviate towards a certain direction throughout the multi-step fitting process, causing their norms to continually increase. Therefore, the optimization speed of the model’s internal representations should be slowed down. Thus, to tackle this problem, we incorporate a regularization loss on the latent representation of PAEs $\mathbf{Z}_{\boldsymbol{\lambda}}$ to stabilize the exploration, which together regularizes the PAEs with the total variation loss $\mathcal{L}_{\mathrm{TV}}$ [45] on the data space that eliminates the frequency discrepancy between PAEs and physically realizable images:
[…]